# A New Census of the Universe’s Entropy

### v1

@@ 4.3 Relativistic Cosmic Rays
To our knowledge, the entropy associated with high-energy cosmic rays has never been assessed. The central issue with evaluating the entropy associated with such population is that the cosmic-ray flux is measured directly only at Earth, and two extrapolations are therefore in order to capture the population in the entire universe:
@@
Cosmic rays with Milky Way-like spectra are likely confined primarily to galaxies with sufficient magnetic fields to retain them over extended periods. Magnetic fields in galactic halos and disks significantly influence the propagation of cosmic rays, allowing them to undergo the diffusion and energy loss processes observed in the Milky Way. Extragalactic magnetic fields are generally weaker ($\sim 10^{-15}-10^{9}$ G), meaning that cosmic rays in intergalactic space are not expected to have the same spectral characteristics due to rapid escape and different energy loss mechanisms (e.g., interactions with the CMB via pion production for high-energy protons) [113].
@@ 5 Gravitational Wave and Dark Radiation Diffuse Backgrounds
$s_{G}=2\frac{2\pi^{2}}{45}T_{G}^{3}\lesssim 1.66\times 10^{7}k_{B}\ \rm{m}^{-3}.$ (5.2)
@@
$S_{G}=2\frac{2\pi^{2}}{45}T_{G}^{3}V_{\rm{obs}}\lesssim 5.84\times 10^{87}\ k_{B}.$ (5.3)
@@
$\rho_{G}=\frac{\pi^{2}}{30}T_{G}^{4};\quad s_{G}=2\frac{2\pi^{2}}{45}T_{G}^{3}=2\frac{2\pi^{2}}{45}\left(\frac{30\rho_{G}}{\pi^{2}}\right)^{3/4}=\frac{4\times 30^{3/4}\pi^{1/2}}{45}(a_{\rm{dec}}\varepsilon\rho_{m})^{3/4},$ (5.5)
@@
We thus conclude that dark radiation could potentially be a very significant contributor to the overall entropy budget of the universe, unlike what was previously claimed in the literature.
@@ 6 Cosmological Dark Matter
$s_{\rm cold\ DM,\ dec}=k_{B}\frac{\rho+P}{T}\simeq\frac{m_{\rm DM}n}{T}\simeq g_{*s,{\rm DM}}(T_{\rm{dec}})\frac{T_{\rm dec}^{3}}{(2\pi)^{3/2}}x_{\rm{dec}}^{5/2}e^{-x_{\rm{dec}}},$ (6.6)
where $x=m_{\rm DM}/T_{\rm dec}\gg 1$, and we assumed that the dark matter’s chemical potential vanishes. Note that, from conservation of entropy in a conformal volume ($\frac{d(sa^{3})}{dt}=0)$), the DM entropy today is redshifted by a factor $(T_{\rm CMB}/T_{\rm dec})^{3}$, thus
$s_{\rm cold\ DM,\ today}=g_{*s,{\rm DM}}(T_{\rm{dec}})\frac{T_{\rm CMB}^{3}}{(2\pi)^{3/2}}x_{\rm dec}^{5/2}e^{-x_{\rm dec}},$ (6.7)
@@
[…]
@@
[…]
We thus conclude that the thermal limits discussed above provide a firm the upper limit to the possible entropy in the cosmological dark matter sector.
@@ 6.1 Effect of clustering in halos at late times
Assuming the cosmological, smooth component of the DM has a thermal distribution resulting from decoupling at a temperature $T_{d}$, we consider here the effect of late-time clustering in halos.
@@
[…]
where $T(M)\simeq 3400\ {\rm eV}(M/10^{14}\ M_{\odot})$. For a cold relic, since $T_{d}>T_{\rm eq}>T_{0}$, the redshifted value of $T_{d}$ is $T_{0}=T_{d}/z_{d}$. Notice that the redshift of decoupling $z_{d}=\frac{T_{d}}{T_{\rm eq}}z_{\rm eq}$, thus $T_{0}=T_{\rm eq}/z_{\rm eq}\simeq 0.012$ eV. As a result, the effect of clustering on the average temperature of the dark matter is $\delta_{c}\simeq 4.2\times 10^{4}$. In turn, this enhances the associated entropy by a factor $S\to S_{c}=S\delta_{c}^{3/2}\simeq 8.7\times 10^{6}\ S$. Notice that this enhancement is an upper bound, as different initial phase space distribution will receive a smaller boost than that of a thermal phase space density.
@@ 7 Black holes of Stellar origin and Supermassive black holes
thus less than 4 orders of magnitude smaller than the SMBH contribution.
@@
$S_{**}=11.2\times 10^{96}\ k_{B}\ (f_{**}=0);$ $\ 4.27\times 10^{96}\ k_{B}\ (f_{**}=1);$ (7.8)
$S_{\rm field}=1.83\times 10^{96}\ k_{B}\ (f_{\rm field}=0.2);$ $\ 4.93\times 10^{96}\ k_{B}\ (f_{\rm field}=1).$ (7.9)
@@ 8.1 Constraints on the abundance of PBHs
At the the largest masses, the relevant constraints are the incredulity limit (lower) and the CMB limit (upper) (see Ref. [155] for a detailed discussion of PBH limits). The CMB constraint is due to the dipole nature of the Cosmic Microwave Background, which would be compromised by the presence of even a single supermassive PBH. The incredulity limit, appearing in the bottom-right corner in our figures, places constraints on how many SMPBHs could possibly exist in a given environment (the observable universe in this case, but this volume can also correspond to a halo or a cluster, implying 1 PBH per halo/cluster). Above a certain mass, it becomes impossible for more than one PBH of said mass to exist in the entire Universe, given the cosmic abundance of matter, making the probability of the existence of this massive of black holes so accidental and rare to be dubbed “incredulous” [160].
@@
Specifically, and in summary from the bottom left, clockwise, in the figure the shaded areas – with the same color conventions as in fig. 10 of Ref. [155], correspond to CMB spectral distortion, disruption of X-ray binaries, dynamical constraints from Eridanus II, Lyaman-alpha forest constraints, dynamical friction, large scale cosmic structure, compact radio sources, CMB dipole, and the incredulity limit.
@@ 8.2 Maximal entropy of Primordial Black Holes
of width $\sigma=2$, and a power-law with $\gamma=1$ and $-1$, respectively. The latter have the functional form
@@
in a range $[M_{\rm min},M_{\rm max}]$. Note that for $\gamma\neq 1$, either the lower or the upper cutoff can be neglected, so the mass function has effectively only two degrees of freedom. The “central value” is defined here as
@@ 8.3 Extrapolation of constraints for super-massive PBHs
We also consider three other constraints which appear in both Fig. 2 and Fig. 3. The first is a constant value extending from the lower right edge of the Planck constraint. The final two assume a form of $f(M)\propto\exp{(-\sqrt{\lambda\log(M)})}$ with $\lambda=0.65$ curving down and intersecting the incredulity limit at $f(M)\approx 10^{-10}$ and $\lambda=50$ curving much lower, intersecting the incredulity limit at $M\approx 2\times 10^{40}$g. With the more restrictive constraints, such as the “constant” extension, not only is the entropy lower at each given mass, but the SMPBHs are also prohibited from reaching higher masses by the incredulity limit, therefore doubly constraining the highest possible entropy. In the case of the “constant” constraint, the highest mass before it intersects the incredulity limit is approximately $M=6\times 10^{47}$g. This is reflected in the graphs on the right of Fig. 2 and Fig. 3, where the purple line ends abruptly at that same mass of $6\times 10^{47}$g. Generally (with the excpetion of the $\lambda=50$ case) the maximal entropy corresponds to the largest allowed value of the mass distribution peak.
The right panels of Fig. 2 and Fig. 3 show the maximal entropy for the constraints assumed in the left panels, compared to the cosmic event horizon entropy (shown by a black line at the top of the panel). The lines terminate when they reach the corresponding maximal mass (typically at the incredulity limit region). We assume for the calculation of the entropy a lognormal mass function (see above) with $\sigma=0.1$. The plot shows that if no constraints arise up to the CMB limit (top right in the left panel), supermassive primordial black holes can contain as much entropy as $S_{\rm SMPBH}\lesssim 10^{117}\ k_{B}$. We discuss this upper limit in detail in the next section, accounting for the impact of various different mass functions.

### v2

@@ 4.3 Relativistic Cosmic Rays
To our knowledge, the entropy associated with high-energy cosmic rays has never been assessed. The central issue with evaluating the entropy associated with high-energy cosmic rays is that the cosmic-ray flux is measured directly only at Earth, and two extrapolations are therefore in order to capture the population in the entire universe:
@@
Cosmic rays with Milky Way-like spectra are likely confined primarily to galaxies with sufficient magnetic fields to retain them over extended periods. Magnetic fields in galactic halos and disks significantly influence the propagation of cosmic rays, allowing them to undergo the diffusion and energy loss processes observed in the Milky Way. Extragalactic magnetic fields are generally weaker ($\sim 10^{-15}-10^{-9}$ G), meaning that cosmic rays in intergalactic space are not expected to have the same spectral characteristics due to rapid escape and different energy loss mechanisms (e.g., interactions with the CMB via pion production for high-energy protons) [87].
@@ 5 Gravitational Wave and Dark Radiation Diffuse Backgrounds
$s_{G}=2\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}T_{G}^{3}\lesssim 1.66\times 10^{7}\ k_{B}\ \rm{m}^{-3}.$ (5.2)
@@
$S_{G}=2\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}T_{G}^{3}V_{\rm{obs}}\lesssim 5.84\times 10^{87}\ k_{B}.$ (5.3)
@@
$\rho_{G}=\frac{\pi^{2}}{30}\frac{k_{B}^{4}}{\hbar^{3}c^{5}}T_{G}^{4};\quad s_{G}=2\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}T_{G}^{3}=2\frac{2\pi^{2}}{45}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}\left(\frac{30\rho_{G}}{\pi^{2}}\right)^{3/4}=\frac{4\times 30^{3/4}\pi^{1/2}}{45}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}(a_{\rm{dec}}\varepsilon\rho_{m})^{3/4},$ (5.5)
@@
We thus conclude that dark radiation could potentially be a very significant contributor to the overall entropy budget of the universe, unlike what was previously claimed in the literature. Note that while our estimate shows that a dark radiation component could, under certain assumptions, carry an entropy comparable to or larger than known radiation backgrounds, it is important to qualify the plausibility of this possibility. Unlike entirely speculative exotic sectors, dark radiation is a feature of many well-motivated extensions to the Standard Model, such as those involving sterile neutrinos, axion-like particles, or hidden $U(1)$ gauge fields. These components are actively constrained by cosmological observations through their contribution to the effective number of relativistic degrees of freedom $\Delta N_{\mathrm{eff}}$ during Big Bang Nucleosynthesis and recombination. Current bounds allow for small but nonzero contributions to $\Delta N_{\mathrm{eff}}$, making dark radiation a moderately speculative but observationally testable possibility.
@@ 6 Cosmological Dark Matter
In greater detail, the behavior of $g_{*}(T)$ above $T\sim 150$ MeV ($t\sim 10^{-6}$ s) involves several key considerations:
  • QCD Hagedorn Temperature ($T_{H}^{\rm QCD}\sim 170$ MeV): In the Standard Model, lattice QCD calculations show $g_{*}$ saturates at $\sim 47.5$ after the quark-hadron transition [107]. No unlimited growth occurs in QCD - the exponential growth of hadronic states terminates at the phase transition to quark-gluon plasma [108].
  • String-Theoretic Hagedorn Temperature ($T_{H}^{\rm string}\sim 10^{30}$): In string theory, the Hagedorn temperature represents a limiting temperature where the thermal partition function diverges [109]. Near $T_{H}^{\rm string}$, the density of string states grows exponentially:
    $\rho(E)\sim e^{\beta_{H}E},\quad\beta_{H}=(k_{B}T_{H}^{\rm string})^{-1}$ (6.6)
    This suggests effective $g_{*}$ could grow without bound in idealized scenarios. However, holographic entropy bounds and finite-$N$ effects ultimately limit physical $g_{*}$ values [110].
  • Empirical Constraints: While speculative models permit large $g_{*}$, observational constraints from:
    • Big Bang nucleosynthesis ($T\sim 1$,MeV) limit $g_{\lesssim}10^{4}$ [111]
    • Gravitational coupling running constrains $g_{*}\lesssim 10^{32}$ [4]
    • Holographic entropy bounds impose $g_{*}\lesssim S_{\rm CEH}/S_{\rm SM}\sim 10^{33}$ [18]
Of course, this does not preclude the addition of additional, secluded degrees of freedom, that may or not contribute to $\Delta N_{\rm eff}$ at times before BBN.
@@
$s_{\rm cold\ DM,\ dec}=k_{B}\frac{\rho+P}{T}\simeq\frac{m_{\rm DM}c^{3}\ n}{T}\simeq g_{*s,{\rm DM}}(T_{\rm{dec}})\frac{T_{\rm dec}^{3}}{(2\pi)^{3/2}}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}x_{\rm{dec}}^{5/2}e^{-x_{\rm{dec}}},$ (6.7)
where $x=m_{\rm DM}/T_{\rm dec}\gg 1$, and we assumed that the dark matter’s chemical potential vanishes see e.g. [112]. Note that, from conservation of entropy in a conformal volume ($\frac{d(sa^{3})}{dt}=0)$), the DM entropy today is redshifted by a factor $(T_{\rm CMB}/T_{\rm dec})^{3}$, thus
$s_{\rm cold\ DM,\ today}=g_{*s,{\rm DM}}(T_{\rm{dec}})\frac{T_{\rm CMB}^{3}}{(2\pi)^{3/2}}\frac{k_{B}^{4}}{\hbar^{3}c^{3}}x_{\rm dec}^{5/2}e^{-x_{\rm dec}},$ (6.8)
@@
$\frac{s[f]}{k_{B}}\equiv\frac{g}{2\pi^{2}}\int_{m}^{\infty}E^{2}\left(f(E)\ln\left(\frac{1}{f(E)}-1\right)-\ln(1-f(E))\right)dE\simeq\frac{g}{2\pi^{2}}\int_{m}^{\infty}E^{2}f\ln fdE;$ (6.14)
@@
$\rho_{\rm DM}=\frac{g}{2\pi^{2}}\int_{m}^{\infty}\sqrt{E^{2}-m^{2}}\ E^{2}f(E)dE.$ (6.15)
We thus conclude that the thermal limits discussed above provide a firm upper limit to the possible entropy in the cosmological dark matter sector.
@@ 6.1 Effect of clustering in halos at late times
Assuming the cosmological, smooth component of the DM has a thermal distribution resulting from decoupling at a temperature $T_{\rm dec}$, we consider here the effect of late-time clustering in halos.
@@
$\int_{\rm All\ halos}\frac{T(m)}{T_{0}}\ m\frac{dN}{dm}dm\simeq\left(\frac{520\ {\rm eV}}{T_{0}}\right),$ (6.16)
where $T(M)\simeq 3400\ {\rm eV}(M/10^{14}\ M_{\odot})$. For a cold relic, since $T_{\rm dec}>T_{\rm eq}>T_{0}$, the redshifted value of $T_{\rm dec}$ is $T_{0}=T_{\rm dec}/z_{d}$. Notice that the redshift of decoupling $z_{d}=\frac{T_{\rm dec}}{T_{\rm eq}}z_{\rm eq}$, thus $T_{0}=T_{\rm eq}/z_{\rm eq}\simeq 0.012$ eV. As a result, the effect of clustering on the average temperature of the dark matter is $\delta_{c}\simeq 4.2\times 10^{4}$. In turn, this enhances the associated entropy by a factor $S\to S_{c}=S\delta_{c}^{3/2}\simeq 8.7\times 10^{6}\ S$. Notice that this enhancement is an upper bound, as different initial phase space distribution will receive a smaller boost than that of a thermal phase space density.
We emphasize that the large entropy values potentially associated with cosmological dark matter—comparable to or exceeding the entropy of the cosmic event horizon—are derived under highly speculative and extreme assumptions, including maximal clustering enhancement and the existence of up to $g_{\ast}\sim 10^{32}$ relativistic degrees of freedom in a hidden dark sector. These configurations should be interpreted as theoretical upper limits rather than realistic scenarios.
Under these extreme assumptions, if present-day matter entropy exceeds 10% of the present CEH entropy, then the matter entropy would exceed the difference between the asymptotic future CEH entropy and the present CEH entropy. In such cases, as matter temperatures redshift to zero in the far future while the CEH entropy approaches its asymptotic value, the total entropy (matter plus horizon) could potentially decrease. However, we note that such scenarios would require an extraordinarily large number of dark sector species ($\sim 10^{39}$), which would fundamentally alter the gravitational physics through renormalization of the gravitational constant. We leave a detailed study of this point to future investigations.
For practical constraints, assuming the GSL holds and that matter entropy should not exceed the growth in CEH entropy from present to asymptotic future, we derive constraints on dark degrees of freedom (assuming no clustering effects):
$g_{\star s,DM}(T_{dec})\left(\frac{x_{dec}}{20}\right)^{5/2}e^{-x_{dec}/20}\lesssim 2.99\times 10^{39}$ (6.17)
We finally note that Cold Dark Matter (CDM) clusters more efficiently than Hot Dark Matter (HDM), and our maximal entropy boost factor of $\sim 10^{6}$ represents an upper bound applicable to CDM. In the case of HDM, the integral over structures is cutoff at large scales, fixed by the free-streaming of the DM, and effectively the clustering boost is of order unity.
@@ 6.2 Topological Defects
6.3 The Low-Entropy Initial State Paradox
The observed homogeneity of the CMB radiation ($\Delta T/T\sim 10^{-5}$) initially appears to represent a high-entropy thermal equilibrium state. However, when gravitational degrees of freedom are properly accounted for [122], this configuration actually corresponds to an extraordinarily low-entropy initial condition. This arises from two key considerations:
  1. 1.
    Gravitational vs. Thermal Entropy: For a self-gravitating system, the maximum entropy state occurs when matter collapses into black holes [9]. The homogeneous matter distribution observed in $\delta\rho/\rho$ inferred from the CMB represents a highly ordered (low gravitational entropy) configuration:
    $S_{\text{grav}}\propto-\int\rho(\mathbf{x})\ln\rho(\mathbf{x})d^{3}x$ (6.18)
    where $\rho(\mathbf{x})$ is the matter density field (see e.g. [123, 124] for derivations and justifications). The near-uniform $\rho(\mathbf{x})$ at recombination, corresponding to the very small values of $\delta\rho/\rho$ inferred from the small temperature fluctuation of the CMB minimizes $S_{\text{grav}}$ because of the logarithmic term in the expression above. Note that this expression for $S_{\rm grav}$ is not the total entropy of the system, but rather a proxy for the configurational entropy associated with the matter distribution under gravity. Unlike thermal entropy, which increases with homogenization and thermal equilibrium, gravitational entropy increases with the growth of inhomogeneities. In self-gravitating systems, the clumping of mass — ultimately into black holes — corresponds to states of higher entropy [125, 126, 123]. Conversely, a nearly uniform mass distribution minimizes $S_{\rm grav}$, as the logarithmic term penalizes spatial homogeneity. In the context of cosmology, the uniform matter field observed at recombination represents a low gravitational entropy configuration. This is consistent with the idea that the early Universe began in a state of extraordinarily low gravitational entropy, despite its high thermal entropy, a concept formalized in Penrose’s Weyl curvature hypothesis [127]. In this view, the vanishing Weyl tensor of the early FLRW spacetime reflects the absence of gravitational clumping, and hence minimal gravitational entropy. For recent discussions connecting these ideas to modern theoretical developments, see e.g. [128].
  2. 2.
    Entropy Partitioning: The total entropy comprises thermal ($S_{\text{thermal}}$) and gravitational ($S_{\text{grav}}$) components:
    $S_{\text{total}}=S_{\text{thermal}}+S_{\text{grav}}$ (6.19)
    While $S_{\text{thermal}}$ was maximal in the early universe’s radiation-dominated plasma, $S_{\text{grav}}$ was minimal due to homogeneity. The subsequent growth of $S_{\text{grav}}$ through structure formation drives the cosmic entropy increase [129].
  3. 3.
    Inflationary Initial Conditions: The observed low gravitational entropy finds natural explanation in cosmic inflation [130], which establishes:
    • Homogeneous spatial curvature ($\nabla\Phi\approx 0$)
    • Scale-invariant density perturbations ($\delta\rho/\rho\sim 10^{-5}$)
    • Initial conditions far from gravitational equilibrium
    These properties create the entropy gradient required for subsequent structure formation.
The CMB’s thermal equilibrium (high $S_{\text{thermal}}$) thus coexists with minimal gravitational entropy ($S_{\text{grav}}$), making the early universe a low-entropy state when all degrees of freedom are considered. This resolves the apparent paradox and aligns with the second law’s requirement for entropy growth over cosmic time.
@@ 7 Black holes of Stellar origin and Supermassive black holes
thus more than 4 orders of magnitude smaller than the SMBH contribution.
@@
$S_{**}=11.2\times 10^{96}\ k_{B}\ (f_{**}=0);\$$\ 4.27\times 10^{96}\ k_{B}\ (f_{**}=1);$ (7.8)
$S_{\rm field}=1.83\times 10^{96}\ k_{B}\ (f_{\rm field}=0.2);\$$\ 4.93\times 10^{96}\ k_{B}\ (f_{\rm field}=1).$ (7.9)
@@ 8.1 Constraints on the abundance of PBHs
At the largest masses, the relevant constraints are the incredulity limit (lower) and the CMB limit (upper) (see Ref. [144] for a detailed discussion of PBH limits). The CMB constraint is due to the dipole nature of the Cosmic Microwave Background, which would be compromised by the presence of even a single supermassive PBH. The incredulity limit, appearing in the bottom-right corner in our figures, places constraints on how many SMPBHs could possibly exist in a given environment (the observable universe in this case, but this volume can also correspond to a halo or a cluster, implying 1 PBH per halo/cluster). Above a certain mass, it becomes impossible for more than one PBH of said mass to exist in the entire Universe, given the cosmic abundance of matter, making the probability of the existence of this massive of black holes so accidental and rare to be dubbed “incredulous” [149].
@@
Specifically, and in summary from the bottom left, clockwise, in the figure the shaded areas – with the same color conventions as in fig. 10 of Ref. [144], correspond to CMB spectral distortion, disruption of X-ray binaries, dynamical constraints from Eridanus II, Lyman-alpha forest constraints, dynamical friction, large scale cosmic structure, compact radio sources, CMB dipole, and the incredulity limit.
8.2 Constraints on the abundance of PBHs
Supermassive PBHs would imprint temperature anisotropies in the CMB through two primary mechanisms:
  1. 1.
    Gravitational Redshift (Sachs-Wolfe Effect): A SMPBH of mass $M$ at distance $d$ creates a gravitational potential $\Phi\sim-GM/(c^{2}d)$, inducing a temperature perturbation:
    $\frac{\Delta T}{T}\sim\frac{|\Phi|}{c^{2}}\approx 1.5\times 10^{-7}\left(\frac{M}{10^{18}M_{\odot}}\right)\left(\frac{d}{10,\text{Gpc}}\right)^{-1}$ (8.1)
  2. 2.
    Dipole from Relative Motion: If an SMPBH moves with velocity $v$ relative to the CMB frame, it generates a Doppler dipole:
    $\frac{\Delta T}{T}\sim\frac{GMv}{c^{3}d}\approx 5\times 10^{-8}\left(\frac{M}{10^{18}M_{\odot}}\right)\left(\frac{v}{600,\text{km/s}}\right)\left(\frac{d}{10,\text{Gpc}}\right)^{-1}$ (8.2)
The observed CMB dipole ($\Delta T/T\sim 10^{-3}$) requires any SMPBH-induced anisotropy to satisfy $\Delta T/T\lesssim 0.1\%$. For SMPBHs within the Hubble volume ($d\sim c/H_{0}$), this implies:
$M_{\text{SMPBH}}\lesssim 10^{18}M_{\odot}\left(\frac{\Delta T/T}{0.1\%}\right)$ (8.3)
This constraint excludes individual SMPBHs more massive than $\sim 1\%$ of the horizon mass while remaining consistent with the observed CMB isotropy.
@@ 8.3 Maximal entropy of Primordial Black Holes
of width $\sigma=2$, and a power-law with $\gamma=1$ and 1-1, respectively. The latter have the functional form
@@
in a range $M_{\rm min},M_{\rm max}]$. Note that for $\gamma\neq 1$, either the lower or the upper cutoff can be neglected, so the mass function has effectively only two degrees of freedom. The “central value” is defined here as
@@ 8.4 Extrapolation of constraints for super-massive PBHs
We also consider three other constraints which appear in both Fig. 2 and Fig. 3. The first is a constant value extending from the lower right edge of the Planck constraint. The final two assume a form of $f(M)\propto\exp{(-\sqrt{\lambda\log(M)})}$ with $\lambda=0.65$ curving down and intersecting the incredulity limit at $f(M)\approx 10^{-10}$ and $\lambda=50$ curving much lower, intersecting the incredulity limit at $M\approx 2\times 10^{40}$g. With the more restrictive constraints, such as the “constant” extension, not only is the entropy lower at each given mass, but the SMPBHs are also prohibited from reaching higher masses by the incredulity limit, therefore doubly constraining the highest possible entropy. In the case of the “constant” constraint, the highest mass before it intersects the incredulity limit is approximately $M=6\times 10^{47}$g. This is reflected in the graphs on the right of Fig. 2 and Fig. 3, where the purple line ends abruptly at that same mass of $6\times 10^{47}$g. Generally (with the exception of the $\lambda=50$ case) the maximal entropy corresponds to the largest allowed value of the mass distribution peak. For constant constraints, this maximal entropy corresponds to $S\sim 10^{104}\ k_{B}$.
The right panels of Fig. 2 and Fig. 3 show the maximal entropy for the constraints assumed in the left panels, compared to the cosmic event horizon entropy (shown by a black line at the top of the panel). The lines terminate when they reach the corresponding maximal mass (typically at the incredulity limit region). We assume for the calculation of the entropy a lognormal mass function (see above) with $\sigma=0.1$. The plot shows that if no constraints arise up to the CMB limit (top right in the left panel), supermassive primordial black holes with masses up to $M_{\rm max}\sim 6\times 10^{54}\ {\rm g}$ can contain as much entropy as $S_{\rm max}\sim 10^{118}\ k_{B}$. We discuss this upper limit in detail in the next section, accounting for the impact of various different mass functions.